# What Can RL Bring to VLA Generalization? An Empirical Study

### v2

@@ A.4 Details of tasks for generalization evaluation
As described in Sec. 5.1, we build multiple tasks to test the generalization of VLA modes. The details of each task is detailed as below:
@@ B.1 PPO performance on generation temperature
We evaluate the performance of PPO with different generation temperatures during training, as shown in Fig. 11. The results show that a very large temperature hinders the training process, while a moderate or small temperature leads to good performance. Therefore, we use a temperature of $1.0$ in PPO.
@@ B.2 PPO performance on LoRA rank
We evaluate the performance of PPO with different LoRA ranks, as shown in Fig. 12. The results show that a lower LoRA rank leads to slightly more efficient training. In all of our fine-tuning experiments, we use a LoRA rank of $32$ to ensure sufficient model capacity. A more detailed study of different LoRA ranks is left for future work.
@@ B.3 Full results of OOD tasks
Detailed results of Fig. 7 are listed in Tabs. 1, 2 and 3, where grasp accuracy, continuous grasp accuracy and task success rate are reported.
Table 1: Detailed results on Vision tasks, with the best results across SFT and RL highlighted in bold. Standard deviations are reported in parentheses.
@@
Table 2: Detailed results on Semantics tasks, with the best results across SFT and RL highlighted in bold. Standard deviations are reported in parentheses.
Acc. Cont. Acc. Suc.
SFT RL SFT RL SFT RL
IND 0.922 (.013) 0.974 (.007) 0.906 (.026) 0.948 (.007) 0.781 (.013) 0.938 (.013)
@@
Table 3: Detailed results on Execution tasks, with the best results across SFT and RL highlighted in bold. Standard deviations are reported in parentheses.
Acc. Cont. Acc. Suc.
SFT RL SFT RL SFT RL
IND 0.922 (.013) 0.974 (.007) 0.906 (.026) 0.948 (.007) 0.781 (.013) 0.938 (.013)

### v3

@@ A.4 Details of tasks for generalization evaluation
As described in Sec.˜5.1, we build multiple tasks to test the generalization of VLA modes. The details of each task is detailed as below:
@@ B.1 PPO performance on generation temperature
We evaluate the performance of PPO with different generation temperatures during training, as shown in Fig.˜10. The results show that a very large temperature hinders the training process, while a moderate or small temperature leads to good performance. Therefore, we use a temperature of $1.0$ in PPO.
@@ B.2 PPO performance on LoRA rank
We evaluate the performance of PPO with different LoRA ranks, as shown in Fig.˜11. The results show that a lower LoRA rank leads to slightly more efficient training. In all of our fine-tuning experiments, we use a LoRA rank of $32$ to ensure sufficient model capacity. A more detailed study of different LoRA ranks is left for future work.
@@ B.3 Full results of OOD tasks
Detailed results of Fig.˜7 are listed in Tab.˜1, where grasp accuracy, continuous grasp accuracy and task success rate are reported.
Table 1: Detailed results across different tasks, with the best results across SFT and RL highlighted in bold. Standard deviations are reported in parentheses.
@@
B.4 Performance comparison on sim-to-real generalization
As discussed, all of our previous experiments are conducted in simulation for scalability and reproducibility when comparing RL and SFT. Nonetheless, real-world performance is a critical factor to validate. We therefore include a preliminary real-world experiment to assess sim-to-real generalizability.
Specifically, we replaced the WidowX arm with a Franka Panda and aligned the camera views. Using the same pipeline as in our main experiments, we retrained VLA models with both RL and SFT under this configuration, and deployed them zero-shot on the real robot. This naturally introduces an OOD setting involving kinematic differences, varied backgrounds, and different real-world objects and receptacles.
In simulation, the training environment contains 16 objects (as described in Sec.˜5.1), 7 backgrounds, and a plate as the receptacle. The real-world evaluation involves 6 new objects (pepper, cucumber, banana, mangosteen, kiwi, sponge), a table as the background, and a green bowl as the receptacle. In both simulation and real-world setups, the initial positions of objects and receptacles are randomized. For real-world evaluation, the initial conditions of each trail are aligned across SFT and RL to ensure a fair comparison. The training and evaluation environments are illustrated in Fig.˜12.
Figure 12: Visualization of the training and evaluation environments for the preliminary real-world evaluation: (a)-(b) examples of simulation training environments; (c)-(d) examples of real-world evaluation settings.
We report the average performance over 30 real-world trials in Tab.˜2. The results show that RL outperforms SFT in both grasp and pick-and-place success rates, demonstrating stronger generalization ability. Qualitatively, we observed that SFT often exhibits excessive movement and overshooting, leading to suboptimal grasp poses and failures. In contrast, RL shows some jitter but successfully iteratively adjusts the end-effector pose, leading to higher grasp success.
Table 2: Preliminary real-world evaluation results. RL fine-tuning in simulation helps sim-to-real transfer compared to SFT.
SFT RL
Grasp Success Rate 0.10 0.43
Pick-and-Place Success Rate 0.00 0.27
B.5 Performance comparison with action chunk
To evaluate the robustness of our conclusions with stronger architectures, we additionally test with OpenVLA-OFT (Kim et al., 2025), which enhance OpenVLA with action chunking.
We adopted an action chunking strategy with a chunk size of 4, enabling the model to predict sequences of actions rather than single steps. To initialize training, we first collected 400 trajectories through motion planning without applying any action filtering, and then trained a warm‑up model using a checkpoint from Huggingface [^4] Haozhan72/Openvla-oft-SFT-libero10-trajall.
Building on this, we conducted RL fine‑tuning in which each observation corresponds to a sequence of 4 actions and is optimized based on cumulative rewards on the following 4 steps. Given the increased dimensionality of the action space, we adjusted the PPO algorithm by introducing action‑dimension‑wise clipping, reducing the clipping ratio from 0.2 to 0.1, and tuning the hyperparameters with discount factor $\gamma=0.96$ and GAE $\lambda=0.85$ to improve training stability.
Results are summarized in Fig.˜13 and Tab.˜3, showing that RL maintains its advantage over SFT even under this more advanced architecture.
Figure 13: Performance comparison between SFT and RL across different tasks with action chunking. Both success rate and relative performance drop are reported.
Table 3: Detailed results across different tasks with action chunking, with the best results across SFT and RL highlighted in bold. Standard deviations are reported in parentheses.
Acc. Cont. Acc. Suc.
SFT RL SFT RL SFT RL
IND 0.984 (.013) 0.969 (.022) 0.953 (.026) 0.953 (.022) 0.776 (.029) 0.891 (.046)
Table 0.964 (.007) 0.948 (.019) 0.953 (.013) 0.948 (.019) 0.812 (.046) 0.885 (.041)
Texture-w 0.844 (.046) 0.885 (.015) 0.807 (.032) 0.870 (.007) 0.547 (.022) 0.792 (.029)
Texture-s 0.719 (.034) 0.781 (.056) 0.661 (.015) 0.734 (.046) 0.401 (.019) 0.578 (.071)
Noise-w 0.932 (.015) 0.974 (.007) 0.880 (.019) 0.958 (.027) 0.703 (.046) 0.870 (.053)
Noise-s 0.755 (.039) 0.839 (.041) 0.724 (.019) 0.786 (.045) 0.401 (.058) 0.625 (.022)
Obj. 0.807 (.007) 0.901 (.019) 0.740 (.007) 0.875 (.022) 0.542 (.037) 0.693 (.037)
Recep. 0.927 (.027) 0.932 (.045) 0.885 (.019) 0.911 (.053) 0.609 (.013) 0.682 (.032)
Instruct 0.620 (.007) 0.641 (.013) 0.604 (.019) 0.630 (.019) 0.479 (.029) 0.562 (.026)
M-Obj. (IND) 0.661 (.075) 0.745 (.019) 0.641 (.080) 0.714 (.032) 0.542 (.059) 0.641 (.066)
M-Obj. (OOD) 0.589 (.048) 0.661 (.019) 0.552 (.041) 0.620 (.019) 0.365 (.032) 0.495 (.048)
Disturb Recep. 0.901 (.027) 0.870 (.029) 0.891 (.022) 0.849 (.048) 0.599 (.039) 0.719 (.056)
M-Recep. 0.891 (.034) 0.859 (.038) 0.849 (.060) 0.828 (.058) 0.365 (.032) 0.297 (.056)
Obj. Pos. 0.807 (.015) 0.917 (.027) 0.771 (.015) 0.896 (.029) 0.453 (.022) 0.682 (.045)
Robot Pose 0.557 (.019) 0.797 (.034) 0.484 (.034) 0.734 (.046) 0.281 (.013) 0.589 (.029)
Obj. Rep. 0.219 (.038) 0.510 (.015) 0.177 (.041) 0.453 (.013) 0.068 (.053) 0.318 (.007)
B.6 Extended performance comparison on opening articulated objects
To move beyond simple pick-and-place and introduce richer manipulation dexterity, we design an open articulated objects task, and replace the WidowX robot arm with a Franka Panda arm to provide a larger working space. The agent must grasp an object’s handle and open the door past $20^{\circ}$. The reward function is shaped as follows: a reward of $0.1$ is given when the handle is first grasped, an additional $0.1$ is granted if the grasp is sustained for five consecutive steps, and a final $1.0$ is provided upon successful door opening.
During training, we introduce randomization across vision, semantics, and execution to improve generalization. For vision, we randomize over 16 background assets (as described in Sec.˜5.1). For semantics, we employ 8 articulated objects that vary in size, name, handle type (bar or boxed), and handle position (up, down, left, or right). For execution, we perturb the object frame with random translations up to $\pm 8\,\text{cm}$ and rotations up to $\pm\pi/48$. The training and evaluation environments are illustrated in Fig.˜14.
We then evaluate generalization on six out-of-distribution (OOD) splits:
  • Vision: 5 novel background scenes,
  • Vision: whole-scene dynamic noise with an image mixing transparency of 0.5,
  • Semantics: 4 unseen articulated objects (microwave, dishwasher, mini-fridge, oven),
  • Semantics: 16 re-phrased language instructions,
  • Execution: translations up to $12\,\text{cm}$ and rotations up to $\pm\pi/24$,
  • Execution: random initial offsets on joints 0/1/2/4 of the robot arm.
As shown in Tab.˜4, RL continues to outperform SFT on semantic and execution splits, consistent with our main pick-and-place results.
Figure 14: Illustration of the open articulated objects task. (a)-(b) Two representative training scenarios. (c) Set of articulated objects used in training (left two rows) and in OOD testing (the most right row).
Table 4: Performance comparison on the open articulated objects task, with results averaged over 3 random seeds. Standard deviations are reported in parentheses.
. SFT RL Success Rate Degradation Ratio Success Rate Degradation Ratio IND 0.745 (.015) 0.724 (.045) New Backgrounds 0.766 (.071) +0.028 0.734 (.058) +0.014 Dynamic Noise 0.526 (.015) -0.294 0.448 (.037) -0.381 Vision Avg. 0.646 (.043) -0.130 0.591 (.048) -0.184 Articulated Obj. 0.042 (.015) -0.944 0.151 (.019) -0.791 Re-phrased Inst. 0.703 (.077) -0.056 0.620 (.041) -0.144 Semantic Avg. 0.373 (.046) -0.500 0.386 (.030) -0.468 Obj. Pose 0.448 (.032) -0.399 0.484 (.089) -0.331 EE Pose 0.151 (.027) -0.797 0.312 (.013) -0.568 Execution Avg. 0.300 (.030) -0.600 0.398 (.051) -0.450